OpenStudy (morgan_home.study):

Use number properties to simplify the following expression. -5 + (5 + 3)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (briellshaw):

-5+8=?

OpenStudy (briellshaw):

8+(-5

OpenStudy (skullpatrol):

recall the associative property of addition

OpenStudy (morgan_home.study):

3?

OpenStudy (skullpatrol):

$$\Huge -5 + (5 + 3) = (-5 + 5) + 3$$

OpenStudy (skullpatrol):

yes, the answer is 3, but the question asks for you to use the properties of real numbers

OpenStudy (skullpatrol):

next use the property of opposites: (-5 + 5 ) +3 = 0 +3 and finally the identity property of addition: 0 +3 = 3

OpenStudy (morgan_home.study):

thx

OpenStudy (skullpatrol):

Thanks for asking :-)